A density bottle contains 104.5gms of a liquid at 34degree centigrade on heating it to 100 degree centigrade the mass of liquid expelled is 4.5gms .Then find the coefficient of apparent expansion of liquid.

Grade:11

3 Answers

chaitnyakishore
10 years ago
Let the density of the liquid in the density bottle be `d` Then volume of the liquid at 34degree centigrade V1=md=104.5d volume of the liquid expelled when it is heated to 100degree centigrade(V2-V1)=4.5d Coefficient of apparent expansion of liquid=[V2-V1]/[V1*(100-34)] =[4.5d]/[104.5d*66] =0.000652458
Vyshnavi
10 years ago
0.0007/degree centigrade
sivakrishna
10 years ago
let the density of the liquid in the density bottle be d. given that mass of liquid m1=104.5gms t1=34 m2=m1+4. =104.5+4.5 =109gms we know that coefficient of apparent expansion =(V2-V1)/V1*(t2-t1) =(m2/d-m1/d)/m1/d*(V2-V1) =(109/d-104.5/d)/104.5/d*(66) =(4.5/d)/6897/d =4.5/6897 =0.00065245759 =0.0007/degree centigrade